OpenStudy (anonymous):

find the value of a if the lines ax+3y=5 and 2x+(a+1)y=6 are perpendicular

OpenStudy (anonymous):

2a+3(a+1)=0 solve for a 2a+3a+3=0 5a=-3 a=-3/5

OpenStudy (anonymous):

a1*a2+b1*b2=0 is the condition for perpendicularity so, a*2+3*(a+1)=0 where a1=a;a2=2;b1=3;b2=(a+1) 2a+3a+3=0 5a+3=0 a=-3/5